OpenStudy (anonymous):

-16=3x-4 solve for x

OpenStudy (anonymous):

Add 4 to both sides, then divide by 3

OpenStudy (triciaal):

isolate the term with the variable to one side and constants to the other side of the equation. whatever you do to one side also do to the other to keep the balance. simplify then divide each side by the coefficient of the variable to get 1 time the variable =

OpenStudy (triciaal):

-16 + 4 = 3x -4 + 4 -12 = 3 x -12/3 = 3 x / 3 x = -4